GRADUATE STUDENT AND NEW PHD VOLUNTEERS

Name (26)  @math.uchicago.edu weeks  course preference
Alan Anders  aanders  8  3b, 3c, 4a
Sundeep Balaji  sundeepb  8
Marius Beceanu  mbeceanu  2-5  Fefferman
Ian Biringer  biringer  1-2, 8?
Michael Broshi  mbroshi  1-7, 8
Anna-Marie Bohmann  bohmann  8
Khalid Bou-Rabee  khalid  8
Mitya Boyarchenko  mitya  8
Tom Church  tchurch  8  geometry
Richard LeRoy Cudney  cudney  8
Matthew Day  mattday  8  geometry
Shawn Drenning  shawn  8  any
Damir Dzhalil Dzhafarov  damir  2-8  1, 2c, 2a, 3b
Spencer Dowdall  sdowdall  3-8  Farb, Bader
Jim Fowler  fowler  8
Michael Geline  geline  1-4  Farb
Bertrand Guillou  bertg  1-4  May, Farb, Law-Feff
Niles Johnson  niles  8
John Lind  lind  8
Justin Malestein  justinm  8  Farb, Fiore, Lawler, ...
Aaron Marcus  amarcus  8  mild pref for 4 or 3
Courtney Morris  clmorris  1-4 or 1-6
Irine Peng  irine  8  Lawler, Farb; Kerr
Andrew Putman  andyp  1-4  any but 1, prefer 3a
Travis Schedler  trasched  2-8  Babai
Ian Shipman  ics  8
Saravanan Thiyagarajan  avan  8  2, 3, and 4
Gabe Kerr  gdkerr  1-6

A few graduating undergraduates are also available to help.

Samuel Altschul salt@uchicago.edu  ?
Shawn Henry  shawnh@uchicago.edu 8 weeks

1. BABAI. discrete mathematics. Eight weeks. 2a. LAWLER. Random walks and Brownian motion. Weeks one and two. 2b. FEFFERMAN. When does a Fourier series converge?. Weeks three – four. 2c. FIORE: Mathematics and music. Weeks five and six. 3a. FARB. Mapping class groups. Weeks one through four. 3b. BADER. Groups of two by two matrices. Weeks five and six. 3c. KERR. Toric geometry. Weeks seven and eight. 4a. MAY. Surfaces, algebras, and TQFT’s. Weeks one through four. 4b. ANGELTVEIT. Products on topological spaces. Weeks five through eight.

5. BABAI. APPRENTICE PROGRAM: Number theory and linear algebra

(see next page)

Participants and numbers of mentors; need people in charge by subgroup 31 or 32 apprentices: 8-10 mentors Geline, Guillou, Kerr, Morris, Putman 22 SESAME (weeks 7-8): 6-7 mentors 8 YSP 7-8: 2-3 mentors 8 YSP 9-10: 2-3 mentors 9 YSP 11-12, others: 2-3 mentors

Course assistance (tentative suggestions in parentheses)

Babai needs four or five people: Schedler (in charge), Dzhafarov, Drenning, one or two more (volunteers are needed). Lawler - Fefferman: check with them; maybe two: Beceanu (Peng)? Farb should have four: (Church, Day, Dowdall, Malestein)? May probably wants four: (Guillou, Johnson, Lind, Bohmann)? Check with instructors; they should all have at least one: Bader, Kerr: Anders Angeltveit: Lind